### v1

# Magnetic texture control in ion-implanted metamaterials

Abstract
We study experimentally the impact of the additive fabrication method on the magnetic properties of Fe+-implanted Pd square artificial spin ice lattices. Our findings show that these lattices undergo magnetic collapse at higher temperatures than their continuous film counterparts. This behavior is attributed to the additive fabrication process, which induces an inhomogeneous Fe concentration within the lattice building blocks. Moreover, the implantation process creates a magnetic depth profile, enabling temperature-dependent tunability of the magnetic thickness. These additional internal degrees of freedom broaden the design possibilities for magnetic metamaterials, allowing precise fine-tuning of their static and dynamic properties to achieve complex and customizable behaviors.
I Introduction
Complex systems are characterized by interactions among their components that lead to the spontaneous emergence of collective properties. In condensed matter physics, magnetic metamaterials have become an exemplary platform for investigating such complexity. These systems consist of arrays of ferromagnetic elements of sub-m size that interact magnetostatically [1, 2], with these interactions giving rise to emergent phenomena, including collective magnetic order [3, 4, 5, 6, 7, 8] and dynamics [9, 10, 11, 12, 13, 14, 15]. The building blocks of magnetic metamaterials, mesoscopic ferromagnetic elements, are often approximated as point-like magnetic dipoles or artificial magnetic atoms. While these analogies are useful, they fall short in capturing phenomena such as thermal fluctuations and phase transitions. Recent studies have provided a more nuanced analysis of the intrinsic magnetic properties and dynamics of these elements [16, 17, 18, 19, 20]. In a notable example, the initially simulated internal degrees of freedom, leading to excess entropy and modification of dynamics as reported by Gliga et al. [16], were later experimentally verified, with their temperature-dependent behavior characterized in detail by Skovdal et al. [21]. These systems bridge theoretical models of complexity with experimental realizations, enabling studies of the interplay between competing interactions [22, 23, 24], frustration, and exotic magnetic phases [25, 26]. In addition to their fundamental significance, the complexity in magnetic metamaterials has fueled technological advancements in fields such as spintronics [27], data storage [28], and unconventional computation [29, 30].
In this work, we explore how magnetic metamaterials can serve as a platform to deepen our understanding of complexity in physics. Using electron-beam lithography and ion implantation techniques [31], we fabricated ferromagnetic elements with intrinsic compositional and magnetic inhomogeneities. Our investigation examines the effects of these inhomogeneities on the ordering temperature of extended arrays and the internal magnetic textures of individual elements. The study focuses on square artificial spin ice (ASI) lattices composed of elongated, stadium-shaped nanoelements—referred to as mesospins. While ASIs have been widely studied for emergent order [3, 5] and thermally induced dynamics [9, 32, 21] in mesospins made from isotropic and homogeneous magnetic materials, our work provides a direct comparison with systems exhibiting intrinsic mesospin inhomogeneities. This comparison not only enhances our understanding of their fundamental behavior but also underscores their potential for practical applications [16], as it enables the deliberate design of both the static and dynamic properties of magnetic metamaterials.
II Materials and Methods
II.1 Sample fabrication
Two samples were studied, having different Fe concentrations. Both were fabricated by implanting 30 keV 56Fe+ into a Pd film through a patterned Cr mask. The Pd film, deposited using DC magnetron sputtering on a MgO substrate, consisted of a 60 nm Pd layer with a 5 nm V adhesion layer and a 6 nm Cr capping layer. This additive fabrication process yields flat FexPd100-x (where stands for at.%) ferromagnetic structures, where Fe has a concentration-depth profile. A detailed description about the fabrication process is provided in Vantaraki et al. [31]. One sample was implanted with a nominal fluence of 2.71016 ions/cm2, while the other was implanted with 31016 ions/cm2. Higher fluences yield higher Fe concentration [33]. In both samples, the stadium-shaped elements have a length of = 470 nm and width = 170 nm, while the edge-to-edge gap between the elements is = 170 nm (see Fig. 1).
II.2 Hysteresis protocols
The temperature-dependent hysteretic behavior of implanted ASIs was studied using a longitudinal magneto-optical Kerr effect (L-MOKE) setup and a vibrating sample magnetometry (VSM) system. In the L-MOKE setup, a -polarized incident laser beam with a wavelength of 659 nm was modulated by a Faraday cell and subsequently reflected by the sample. The reflected beam was passed through an analyzer (with extinction ratio of 105) and then measured using a biased Si photodiode. The detector was connected with a pre-amplifier and lock-in amplifier. A sinusoidal external magnetic field with an amplitude of 50 mT and a frequency of 0.1 Hz was applied along the [11] and [01] directions of the ASI lattice. The 659 nm laser beam achieves a penetration depth of approximately 11.9 nm in Pd film (see Supplemental Material), detecting the magnetization response from the surface of the sample. The cryostat in the L-MOKE system operates up to 400 K. In addition to the MOKE measurements, magnetic hysteresis loops were recorded by Vibrating Sample Magnetometry (VSM) with a 2 T Lakeshore 8600 magnetometer, equipped with a single-stage variable temperature unit that can reach a temperature of 950 K. This technique records the magnetic response from the entire sample volume.
II.3 Magnetic imaging
The magnetic state of the implanted mesospins was directly imaged using photoemission electron microscopy, employing x-ray magnetic circular dichroism (PEEM-XMCD). These experiments were conducted at the CIRCE (bl24) beamline of the ALBA synchrotron [34]. The images were acquired at room temperature with an x-ray photon energy tuned to the Fe L3-edge.
II.4 Monte-Carlo simulations
The Fe distribution within the mesospins was investigated by performing Stopping and Range of Ions in Matter (SRIM) simulations [35]. To imitate the experimental conditions, the simulations were performed for 30 keV Fe+ ions implanted in a 600 Å Pd film through a 60 Å Cr capping layer. The angle of ion incidence was set to zero degrees (normal to the surface).
II.5 Micromagnetic simulations
The magnetic texture of the inhomogeneous mesospins was explored by performing micromagnetic simulations using MuMax3 [36]. To mimic the experimental conditions, the simulations were performed with stadium-shaped elements of dimensions 470 170 nm2 arranged on a square lattice. The edge-to-edge gap between the elements was = 170 nm. The elements were considered to be composed of Fe with a thickness 3.18 nm. This thickness corresponds to a continuous Fe layer implanted with a nominal fluence of 2.71016 ions/cm2 (see Supplemental Material). The elements were discretized into grid cells with cell size smaller than the exchange length and equal to 2.5 nm.
III Results and Discussion
We begin by comparing the amount of Fe implanted into the ASI array with that in its reference continuous film that was prepared and implanted next to the lattice, with a nominal fluence of 2.71016 ions/cm2. To do so, we examine the ordering temperature for these two samples. The ordering temperature is defined as the temperature where the spontaneous magnetization vanishes. To a good approximation, this temperature coincides with the temperature where the remanent magnetization disappears [37]. As such, we examine the temperature-dependent behavior of the remanent magnetization M from the MOKE measurements. As shown in Fig. 2 (a), the remanent magnetization vanishes at higher temperatures for the implanted ASIs compared to the reference implanted film, reflecting a higher ordering temperature. It is worth noting that the collapse of the magnetic order occurs at lower temperature for the continuous film compared to the lattice even when considering the saturation magnetization (see Supplemental Material). This observation highlights a higher Fe concentration for the patterned film. As shown schematically in the Fig. 2 (b), the additional Fe concentration in ASIs is attributed to the Fe ions that impinge the mask near the edges. Due to multiple scattering events, some ions manage to scatter back out of the mask, and then become embedded in the Pd film at an angle. As a consequence, the core area of the elements is enriched by additional Fe dopants, that are not implanted in the periphery of the elements or the reference continuous film. Similar mask proximity effects have been reported for CMOS devices [38, 39, 40]. Based on the increase of the critical temperature, we estimate that the ASI lattice has approximately 1-3% higher Fe concentration compared to the continuous film. This estimation was derived from the relationship between the Fe concentration and the ordering temperature found in Refs. [41, 33, 42].
To explore the implanted-Fe distribution, we performed Monte Carlo simulations using the SRIM software. The simulations provide qualitative information about the Fe+ spread along the depth of the Pd film as well as perpendicular to it for a single point in space, without accounting mask proximity effects. As shown in Fig. 1 (d), the simulation results confirm the experimental observations that Fe has a concentration depth-profile [31]. In addition, the calculations demonstrate that Fe exhibits a concentration lateral-profile as a consequence of the lateral straggling of ions. The lateral implantation profile is often neglected in continuous films, as the lateral motion of ions creates a broader distribution due to the system’s extension. However, the concentration lateral-profile can no longer be neglected in the lattices due to the clear boundary between the implanted and unimplanted regions [43]. Taking into account the mask proximity effects, the lateral implanted profile within the lattice building blocks is expected to be further amplified. As such, the implanted mesospins are both vertically and laterally inhomogeneous as a consequence of the ions straggling and mask proximity effects.
To directly illustrate the impact of the inhomogeneities on the magnetic state of mesospins, we image the ASI array at the remanent state by performing XMCD-PEEM experiments. Representative images are shown in Fig. 3 (a) and (b), when the field was applied along the [11] and [01] direction, respectively [44]. In both images, the elements are displayed in white and black contrast, confirming their single domain state. The ASI array has obtained a ferromagnetic (or Type II) remanent state when the field is applied along the [11] direction. Similarly, by applying a field along the [01] direction and then returning to the remanent state, results in a full ferromagnetic order for the elements parallel to the [01] direction. For the mesospins parallel to the [10] direction, a mixture of ferromagnetic and antiferromagnetic configurations is observed. However, the ferromagnetic order is more dominant than the antiferromagnetic one as it creates vertices of lower energy (E < E) (see Supplemental Material). Similar remanent states have been predicted for conventional lithographic ASI lattices [9].
To better resolve whether inhomogeneities influence the mesospin magnetization texture, we proceed by examining the remanent magnetization normalized by the saturation magnetization, Mrem/Msat, from the L-MOKE measurements. This quantity separates the lattice from the material contributions, as Msat is primarily an intrinsic material property [21]. For collinear magnetic texture, Mrem/Msat has been predicted and measured to be equal with 0.707 for the [11] direction, and 0.5 for the [01] direction [9]. However, as shown in Fig. 3 (c), the implanted ASIs exhibit higher Mrem/Msat values at temperatures below 350 K, approximately 0.8 for the [11] direction, and 0.6 for the [01] direction. These higher values reflect the presence of a non-collinear magnetization component. It is important to mention that such non-collinearities in mesospin texture have also been reported for conventional -doped Pd(Fe) ASI lattices at low temperatures [21]. However, the inhomogeneities inherent to implanted patterns seem to amplify the effect further.
Consequently, we perform micromagnetic simulations to investigate the non-collinearities in implanted mesospins. The simulations confirm the presence of a non-collinear magnetization component at the mesospin edges for any periphery lateral thickness (see Supplemental Material). Subsequently, we calculate the ratio Mrem/Msat as a function of the angle between the collinear and non-collinear components, by assuming constant periphery thickness and non-collinearities at the whole mesospin edges (semicircles). For the calculation of the ratio Mrem/Msat the system was not allowed to relax. The results are presented in Fig. 4. For collinear magnetic texture ( = 0), the ratio Mrem/Msat is equal to 0.707 for the [11] direction and 0.5 for the [01] direction, as expected. By introducing non-collinearities ( > 0), Mrem/Msat increases.
To explore the extent of the mesospin inhomogeneity effects, we choose to investigate the same ASI lattice implanted with a higher nominal fluence of 31016 ions/cm2. The pattern was investigated in the [11] direction and measured with the L-MOKE system at the temperature range of 200 - 370 K and the VSM system at the temperature range of 200 - 700 K. The remanent magnetization as a function of the temperature for the ASI lattice and its reference continuous film extracted from the VSM measurements is presented in Fig. 5 (a). The array magnetization collapses at higher temperatures compared to the film magnetization, verifying that the ordering temperature is higher for the array. The temperature-dependent ratio Mrem/Msat from both L-MOKE and VSM measurements is shown in Fig. 5 (b). The MOKE measurements show that Mrem/Msat has a plateau with a height of approximately 0.8, hinting towards non-collinearities in mesospin magnetic texture. However, the corresponding VSM measurements indicate lower values. As MOKE detects the surface magnetization, whereas VSM quantifies the volume magnetization, and the Fe concentration depth-profile decays from the sample surface towards the substrate [Fig. 5 (c)], the difference in Mrem/Msat indicates that the magnetic moments are more rigid close to the sample surface but fluctuate deeper. This observation highlights a tunable magnetic thickness with temperature, for the implanted building blocks.
In conclusion, we have demonstrated that the magnetic order in Fe-implanted artificial spin ice (ASI) lattices collapses at higher temperatures than in their reference continuous films, with the mesospins exhibiting magnetic inhomogeneity both laterally and vertically. Furthermore, the implanted building blocks display a tunable magnetic thickness with temperature, introducing an additional degree of freedom. This tunability may allow the mesospins to adjust their magnetic dimensionality, transitioning from 2D behavior at higher temperatures to 3D behavior at lower temperatures, and enabling modifications to the dipolar coupling between building blocks. These findings open new pathways for tuning the static and dynamic properties of magnetic materials by leveraging fabrication processes that go beyond traditional subtractive methods such as electron-beam lithography and milling. Such advances may significantly impact future applications of magnetic metamaterials. It was recently demonstrated that the ability to design and control magnetic textures in systems such as ASI enables the realization of tailored magnonic regimes. These emerging textures play a crucial role in defining the accessible ground states and governing transitions between magnetization states, thereby shaping the system’s overall dynamics [19]. The engineering of energy landscapes through fabrication techniques like the ion implantation presented here—harnessing the internal degrees of freedom in mesospins—is expected to be important in the advancement of reconfigurable magnetic metamaterials [45], where dynamic behavior simultaneously functions as the readout. Furthermore, these enhanced internal dynamics, combined with the ability to tune mesospin couplings, open promising avenues for the development of reconfigurable reservoir computing networks based on magnetic metamaterials [46, 47, 48]. The thermal behavior described here is also critical for practical implementations of Ising machines using magnetic metamaterials [49]. The design and precise control of their stochastic dynamics are essential prerequisites for these systems to be effectively employed in solving optimization problems [50].
Acknowledgements.
The authors would like to thank Johan Oscarsson and Mauricio Sortica at the Uppsala Tandem Laboratory for help with ion implantations. The authors are thankful for an infrastructure grant by VR-RFI (grant number 2019-00191) supporting the accelerator operation. The authors also acknowledge support from the Swedish Research Council (projects no. 2019-03581 and 2023-06359). CV gratefully acknowledges financial support from the Colonias-Jansson Foundation, Thelin-Gertrud Foundation, Liljewalch and Sederholm Foundation. KI and UBA acknowledge funding from the Icelandic Research Fund project 2410333. We acknowledge Myfab Uppsala for providing facilities and experimental support. Myfab is funded by the Swedish Research Council (2020-00207) as a national research infrastructure. The PEEM-XMCD experiments were performed at CIRCE (bl24) beamline at ALBA Synchrotron with the collaboration of ALBA staff. MF acknowledges support from MICIN through grant number PID2021-122980OB-C54. The authors are deeply grateful for the support from ReMade@ARI, funded by the European Union as part of the Horizon Europe call HORIZON-INFRA-2021-SERV-01 under grant agreement number 101058414 and co-funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10039728) and by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number 22.00187. VK and CV would like to thank Prof. Björgvin Hjörvarsson for discussions.The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Supplemental Material: Magnetic texture control in ion-implanted metamaterials
IV Laser penetration depth on Pd
We estimate the penetration depth of a laser light with wavelength = 659 nm on a Pd material. The penetration depth is the reciprocal of the absorption coefficient :
The absorption coefficient determines how far light can penetrate into a material before it is absorbed and is expressed as:
where is the extinction coefficient. For Pd material and light of wavelength = 659 nm, = 4.42 [51]. Consequently, the absorption coefficient is = 0.084 nm-1 and the penetration depth is = 11.9 nm.
V Thickness of implanted Fe
To estimate the thickness of Fe layer implanted with nominal fluence = 2.71016 ions/cm2, we consider that the implanted Fe has formed a compacted layer. To a first approximation, the mass of a single Fe+ ion is the same with the mass of a single Fe atom and equal to:
where is the atomic mass of Fe and is the Avogadro’s number. Consequently, the area density , i.e total mass of Fe implanted per unit area, is:
By considering that the density of Fe is = 7.87 g/cm3, the thickness of the implanted Fe as compacted layer is:
VI Collapse of the magnetic order
To verify that the ordering temperature is higher for the ASI lattice compared to its reference continuous film that was prepared and implanted next to the lattice with 30 keV Fe of nominal fluence 2.71016 ions/cm2, we also examine the saturation magnetization as a function of the temperature. As shown in Fig. S1, the saturation magnetization vanishes at higher temperature for the lattice compared to the continuous film, reflecting a higher ordering temperature for the ASI array.
VII Vertex statistics for the remanent state [01]
To quantify the magnetic order established within the ASI lattice at the remanence state after the application of a field along the [01] direction, we determine the population of various types of vertex configuration. A vertex at which four elements converge can exhibit four distinct topologies, classified as Type I, II, III, and IV. Type I possesses the lowest energy, and Types II, III, and IV exhibit progressively higher energy. Vertex statistics are presented in Fig. S2. We notice that the ASI lattice has obtained a mixture of Type II and III configurations.
VIII Non-collinearities at mesospin edges
We performed micromagnetic simulations to explore the non-collinearities in mesospin texture. For the simulations, we investigated a vertex using periodic boundary conditions (Fig. S3), while each element is made by a core and a periphery area. These two areas were described by different magnetic properties, saturation magnetization and exchange stiffness constant. Specifically, the saturation magnetization and the exchange stiffness constant for the core area of the elements were chosen based on previous studies (M = 3.5105 A/m and A = 3.3610-12 J/m) [52]. For the periphery, we assumed that the Fe concentration is lower by 2% compared to the core area of the elements, and that the magnetic properties follow a linear relationship with the Fe concentration. Consequently, the saturation magnetization and the exchange stiffness constant for the periphery were considered to be equal with M = 3.43105 A/m and A = 3.292810-12 J/m respectively. The magnetization of each mesospin was assumed uniform, and subsequently the system was relaxed into its ground state. The simulations were performed for various periphery lateral thicknesses. Fig. S3 shows a representative relaxed vertex state, confirming the presence of non-collinear magnetization component at mesospin edges.